# Efficient Transformer-based Hyper-parameter Optimization for Resource-constrained IoT Environments

### v1

@@ V Results and Discussion
This section reports the results of the experimental procedure and analyzes them based on the defined performance metrics and layer-generation interpretability. TRL-HPO results are obtained in the exploitation phase of the offline RL.
@@ Methods
Table II shows the results of TRL-HPO versus SOTA methods that built CNN models from scratch. Based on evaluation results, MetaQNN produces the best model when classifying MNIST data reported in the exploration phase. This result is attributed to the freedom endowed to the controller to generate models up to 12 layers and the integration of priors via the addition of dropout layers every 2 layers [5]. The accumulation of more layers improves classification results; however, at the expense of wider search space that increases the running time. This is evident with the computational footprint of MetaQNN that exceeds TRL-HPO, despite running on 8$\times$ the amount of GPU resources. On the other hand, the approach in [8] is the first to converge based on the reported results, given it is a Multi-agent RL algorithm. However, this method experimented with four configurations, highlighting the limited exploration of this strategy and the pre-mature convergence of the master agent. This limitation is reflected by its poor best models compared to the other two approaches, excluding it from further analysis. Since the larger search space of MetaQNN would not provide a fair comparison with TRL-HPO, it was imperative to compare these two approaches using AccTime. Within TRL-HPO convergence time, its generated models exhibit the best performance compared to the SOTA models. Despite TRL-HPO’s need to generate more models given the requirement of progressive reward, it outperforms other models in the AccTime metric. Compared to other SOTA methods, TRL-HPO can progressively generate CNN models with satisfactory performance in a shorter period. These two conditions are important for resource-constrained IoT environments. In terms of AccTime, on average, TRL-HPO outperforms MetaQNN by 6.8 %.
@@ V-B Layer Analysis
Figure 2 summarizes the distribution of layer combinations with negative rewards. Two combinations stand out, Class 1 represented by Conv2D and Conv2D combinations and Class 2 represented by two consecutive FCLs. A layer combination refers to the layer that produced a negative reward and its previous layer, shedding light on the combinations that degrade model performance. Two main insights can be gathered from Figure 2. The first is that the negative reward is overwhelmingly attributed to the accumulation of FCLs. This means that the stacking of multiple of these layers is superfluous on the MNIST dataset, a conclusion that aligns with the best models obtained in [5]. In a similar vein, the second insight highlights that additional Conv2D layers can be detrimental to model performance, especially with MNIST data of limited visual complexity. As such, these observations are beneficial to garner knowledge about superfluous stacking of layers, suitable for environments with resource constraints.
[…]
[scale=0.4]diff_attn.png
@@
Figure 3 depicts the distribution of the differences in attention values between layers with the best performances and layers with negative rewards. The distribution demonstrates the close fidelity of the attention mechanism and the utility of a specific layer in its classification performance. As such, the attention mechanism captures the close affinity between layers assigning higher weights to promising layers while shunning layers with poor performances. The difference in attention values averages 0.22, a significant value considering that some attention is assigned to the layers that are yet to be generated. In these cases, the attention values of the layer degrading model performances are higher than these to-be-generated layers. Therefore, we conclude that the attention mechanism reflects the relationship between layers within the context of generating best-performing models.
VI Open Challenges and Future Recommedations
The implementation of the TRL-HPO process opens new avenues for exploring the improvement of this process, which spans many research and practical questions about RL-based HPO implementations.
@@ VI-A Computational Time
Despite improvements in the efficiency of the TRL-HPO process, it has yet to converge in a short time. The bottleneck in the training process originates from two main sources: (1) the training of individual models, and (2) optimization rounds of the actor-critic approach. The first source is a necessary evil because of the exploration requirements and the need to evaluate the actor’s actions. Determining the training data size, the number of epochs, and learning rates is required so that close to optimal accuracy is achieved. During the exploration and action assessment phases, repetitive models can be generated. Storing these models is a method to avoid retraining them, which can overwhelm the available GPU or RAM resources, especially with a large ER buffer. An alternative method is storing already trained models in a database or a hash table, whereby the model is represented via its HP or its hash value. A salient issue relates to querying algorithms that should search for HP representations available in the database. The querying time will increase with the expansion of the database, requiring a more intricate search procedure. These methods should be investigated to budget the computational time towards more fruitful procedures that benefit RL-based solutions.
@@ VI-B Exploration
RL implementations depend on the trial-and-error procedure taking place in the explorations phase. However, a shortcoming of this phase, either through DDPG’s random noise or $\epsilon$-greedy algorithms [5], relates to its lack of direction. This means that the agent continues exploring unpromising areas of the search space with relatively poor performances. Therefore, to reduce this unnecessary exploration, it is imperative to include priors or rectify the exploration stage. As such, aspects of the BO process should be integrated into the RL-based strategies to streamline the exploration process. The randomness can be incorporated using a stochastic policy, which is part of the soft actor-critic policy gradient approach [15]. In the exploration stage, the stochastic policy outputs the mean and standard deviation of a normal distribution. The standard deviation is progressively reduced to achieve a deterministic policy based on the reward function. As such, this suggestion opens new frontiers toward examining methods that can associate the uncertainty of the stochastic policy with steering the RL’s agent exploration.
@@ VI-C Reward Function
The progressive reward function devised in TRL-HPO formulation is foundational to gaining insights into the contributions of each layer to performance enhancement. However, this layer accumulation process exposes two limitations. The first limitation pertains to the training of many models, even with a single layer, to infer this performance. This pre-requisite restricts expanding the model generation to a few layers, which applies to our experimental procedure. These constraints can be sidestepped via block generation at the expense of sacrificing transparency and layer diversity. A block is defined as a set of similar basic layers. This small modification is reflected by altering the actor’s output to produce the number of these similar blocks. The second limitation of this process relates to the training procedure. The HPO procedure is conducted on a subset of the training set to facilitate the training and validation procedures. This means that the performance of the generated CNN models can significantly vary, depending on the train and test splits. In light of the stopping criteria, this factor is crucial in deciding whether to halt or continue the layer generation process.
The reward function that drives the model generation process can be adjusted to achieve better results. Instead of performance differences, the reward can be the inverse distance to a perfect metric (100% accuracy in our example). In this way, the rewards are less sparse but require the tuning of a discount rate. Additionally, other metrics, such as f1-score and top-3 accuracy can replace accumulated accuracy. The adoption of any of these metrics depends on the practitioners’ priorities.
VI-D Transferability
With each new contribution toward the HPO procedure, questions of generalizability and transferability loom to undermine their usefulness, which also applies to RL-based solutions. In TRL-HPO, the actor and critic are constructed to generate and evaluate models using basic layers of FCL, CNN, and MaxPool layers, necessary to build DNN models. The transferability comes into question when TRL-HPO trained on one dataset needs to be applied to other datasets. In that regard, the main concern relates to the size of the input sequence, trained on the initial dataset, which constrains the size of the produced models. This raises the question of how to expand the input sequence to involve more layers?.
The answer to this question requires re-defining the concept of the input sequence (state) supplied to the actor and critic. In its current form, each index in the state sequence combines the layer definition and its performance. To enable transferability and the expansion of TRL-HPO to more profound use cases, the actor should output an additional value representing the number of similar layers. The remaining alteration is to determine the upper bound of the permitted number of these layers so that the model’s capacity would match the dataset’s complexity.
@@ VII Conclusion
The HPO process is a fundamental step in the ML pipeline that contributes to enhancing model performance. However, its computational footprint is prohibitive for widespread adoption and current methods overlook the transparency in the layer generation process. These factors are imperative to realize IoT applications’ function, AVs in particular, requiring model partitioning to fulfill the resource constraints of edge environments. To address these limitations, this paper proposes TRL-HPO framework that combines transformers with an RL actor-critic approach. The attention mechanism, parallelization, and the progressive generation of layers are all novel properties of this framework within the transparency and time requirements. These advantageous factors were empirically established compared to SOTA approaches using the MNIST dataset. Moreover, a list of research questions scrutinizing RL-based solutions, including TRL-HPO, is presented along with a set of recommendations. Future work will target the presented questions to design a general-purpose HPO process.

### v2

@@ V Results and Discussion
This section reports and analyzes the results based on the defined performance metrics and layer-generation interpretability. TRL-HPO results are obtained in the exploitation phase of the offline RL.
@@ Methods
Table II shows the results of TRL-HPO versus SOTA methods that built CNN models from scratch. Based on evaluation results, MetaQNN produces the best model when classifying MNIST data reported in the exploration phase. This result is due to the controller’s freedom to generate models up to 12 layers and the integration of priors by adding dropout layers [6]. The accumulation of more layers improves classification results; however, at the expense of wider search space that increases the running time. This is evident with the computational footprint of MetaQNN that exceeds TRL-HPO, despite running on 8$\times$ the amount of GPU resources. On the other hand, the approach in [9] is the first to converge based on the reported results. However, this method experimented with four configurations, highlighting the limited exploration of this strategy and the pre-mature convergence of the master agent. This limitation is reflected by its poor best models compared to the other two approaches, excluding it from further analysis. Since the larger search space of MetaQNN would not provide a fair comparison with TRL-HPO, it was imperative to compare these two approaches using AccTime. Within TRL-HPO convergence time, its generated models exhibit the best performance compared to the SOTA models. Despite TRL-HPO’s need to generate more models given the requirement of progressive reward, it outperforms other models in the AccTime metric. Compared to other SOTA methods, TRL-HPO can progressively generate CNN models with satisfactory performance in a shorter period. These two conditions are important for resource-constrained IoT environments. In terms of AccTime, on average, TRL-HPO outperforms MetaQNN by 6.8 % in accuracy. The improvement in the accuracy result of TRL-HPO compared to MetaQNN is statistically significant (p-value = 0.02) by applying a t-test on the distribution of TRL-HPO versus the 1-sd values of MetaQNN.
In a resource-constrained IoT environment, analyzing the energy consumption and processing power of TRL-HPO compared to SOTA methods is imperative. We analyze these methods using the method’s number of parameters and the floating point operations (FLOPs), which cover IoT-related concerns. The MetaQNN requires storing all models and state transitions to find the best-performing models, which with many combinations of layers and HPs, is prohibitive for the IoT environment. In Neary et al. approach [9], an RL agent is assigned to every HPs and a master agent finds the best combination of individual HPs. For Neary et al., 399K FLOPs are required for inference whereas TRL-HPO requires 991k FLOPs. On the other hand, both models have the same number of model parameters (78.8k parameters).
@@ V-B Layer Analysis
Figure 2 summarizes the distribution of layer combinations with negative rewards. Two combinations stand out, Class 1 represented by Conv2D and Conv2D combinations and Class 2 represented by two consecutive FCLs. A layer combination refers to the layer that produced a negative reward and its previous layer. Two main insights can be gathered from Figure 2. The first is that the negative reward is overwhelmingly attributed to the accumulation of FCLs. This means that the stacking of multiple FCL layers is superfluous on the MNIST dataset, a conclusion that aligns with the best models obtained in [6]. The second insight highlights that additional Conv2D layers can degrade model performance, especially with MNIST data of limited visual complexity. As such, these observations are beneficial to garner knowledge about stacking of layers, suitable for environments with resource constraints.
@@
Figure 3 depicts the distribution of the differences in attention values between layers with the best performances and layers with negative rewards. The distribution demonstrates the close fidelity of the attention mechanism and the utility of a specific layer in its classification performance. As such, the attention mechanism captures the close affinity between layers assigning higher weights to promising layers while shunning layers with poor performances. The difference in attention values averages 0.22, a significant value considering that some attention is assigned to the layers yet to be generated. In these cases, the attention values of the layer degrading model performances are higher than these to-be-generated layers. Therefore, we conclude that the attention mechanism reflects the relationship between layers within the context of generating best-performing models.
Challenge Summary Solutions
Computational Time Training of individual models Optimization rounds of actor-critic approach Storing the trained models in a hash table to avoid re-training
Exploration Strategy DDPG’s lack of direction Application of RL methods with inherent randomness such as soft actor-critic approach
Reward Function Training many models to infer their contribution to classification results Prioritization of accuracy Generating a block of similar layers Inclusion of inference time, energy consumption, and number of operations per second
Reward Function Training many models to infer their contribution to classification results Prioritization of accuracy Generating a block of similar layers Inclusion of inference time, energy consumption, and number of operations per second
Generalizability, Scalability, and Transferability Scalability and Transferability concerns with larger datasets and complex architectures. Generalizability to different IoT Environments Expansion of the input sequence and actor’s block generation Changing the mappings of the actor’s output to adapt to different DNN models
Integration into Automated ML N/A Replacing the HPO process for IoT deployment frameworks
TABLE III: Summary of Challenges
VI Open Challenges and Future Recommendations
The implementation of the TRL-HPO process opens new avenues for exploring the improvement of this process, which spans many research and practical questions about RL-based HPO implementations, summarized in Table III.
@@ VI-A Computational Time
Despite improvements in the efficiency of the TRL-HPO process, it has yet to converge in a short time. The bottleneck in the training process originates from two main sources: (1) the training of individual models, and (2) optimization rounds of the RL approach. The first source is a necessary evil because of the exploration requirements and the need to evaluate the actor’s actions. Determining the training data size, the number of epochs, and learning rates are required to achieve close to optimal accuracy. During the exploration and action assessment phases, repetitive models can be generated. Storing these models can avoid retraining them, which can overwhelm the available GPU or RAM resources, especially with a large ER buffer. An alternative method is storing already trained models in a database or a hash table, whereby the model is represented via its HP or its hash value. A salient issue relates to querying algorithms that should search for HP representations available in the database. The querying time will increase with the expansion of the database, requiring a more intricate search procedure. These methods should be investigated to budget the computational time towards more fruitful procedures that benefit RL-based solutions.
@@ VI-B Exploration
RL implementations depend on the trial-and-error procedure in the exploration phase. However, a shortcoming of this phase, either through DDPG’s random noise or $\epsilon$-greedy algorithms [6], relates to its lack of direction. This means that the agent continues exploring unpromising areas of the search space with relatively poor performances. Therefore, it is imperative to include priors or rectify the exploration stage to reduce unnecessary exploration. Aspects of the BO process should be integrated into the RL-based strategies to streamline the exploration process. The randomness can be incorporated using a stochastic policy, which is part of the soft actor-critic policy gradient approach [15]. In the exploration stage, the stochastic policy outputs the normal distribution’s mean and standard deviation. The standard deviation is progressively reduced to achieve a deterministic policy based on the reward function. This suggestion opens new frontiers toward examining methods that can associate the uncertainty of the stochastic policy with steering the RL’s agent exploration.
@@ VI-C Reward Function
The progressive reward function of TRL-HPO is foundational to gaining insights into the contribution of each layer to performance enhancement. However, this layer accumulation process requires training many models to infer their performance, restricting the model generation process to a few layers. These constraints can be sidestepped by generating similar layers at once, which is referred to as a block generation procedure. As such, computational time is gained at the expense of transparency and layer diversity. In resource-constrained environments, factors such as energy consumption, inference time, and number of computational operations per second can gain precedence over accuracy. These factors can be added to the reward function and assigned a weight based on application requirements.
VI-D Generalizability, Scalability, and Transferability
With each contribution toward the HPO procedure, questions of generalizability, scalability and transferability loom to undermine their usefulness; a concern that applies to RL-based solutions. In TRL-HPO, the actor and the critic generate and evaluate models using FCL, CNN, and MaxPool layers, necessary to build DNN models. When applied to complex datasets and architectures, skip and residual connections can be integrated into the TRL-HPO architecture. Towards that end, two main approaches can be followed: (1) expand the input space to include more layers, which is reflected in the sequence length variable, or (2) change the definition of the actor’s output from the layer and its HPs to include the number of similar layers and their HPs. Adopting any of these approaches should consider the trade-off between the availability of computational resources and the layer generation process’s transparency. The same adaptation procedure can be followed when confronting a transferability concern.
The IoT environment is integrated into various fields, including smart cities, healthcare, buildings, and electric grids. This environment suffers from computational and communication resource scarcity, which can be detrimental to ML applications, such as anomaly detection, object detection, and forecasting. In connection to the IoT environment limitations, the TRL-HPO framework offers solutions manifested in three aspects: (1) its generalizability to different types of DNN models, including CNN, Long short-term memory (LSTM) and FFNN by changing the mappings of the actor’s output, (2) its low computational, storage, and processing footprint, compared to SOTA approaches, and (3) its transparency that enables the reliance on part and not all of the CNN models and exchanging model parameters when the need arises.
VI-E Integration into AutoML
The HPO process is an important procedure in the AutoML pipeline and its enhancement is central to the wide-scale deployment and adoption of these pipelines. The TRL-HPO process is envisioned to replace the generic HPO processes that are part of the ubiquitous cloud computing ML deployment modules. These modules enhance the efficiency and accessibility of ML deployment, by providing user-friendly deployment strategies. TRL-HPO that outperformed SOTA approaches in its convergence time while maintaining good accuracy results benefits experts who have budgetary constraints and desire to prove the viability of their products/models for customers/investors alike.
@@ VII Conclusion
The HPO process is a fundamental step in the ML pipeline that enhances model performance. However, its computational footprint is prohibitive for widespread adoption and current methods overlook the transparency in the layer generation process. These factors are imperative to realize IoT applications’ function, AVs in particular, requiring model partitioning to fulfill the resource constraints of edge environments. To address these limitations, this paper proposes TRL-HPO framework that combines transformers with an RL actor-critic approach. The attention mechanism, parallelization, and the progressive generation of layers are all novel properties of this framework within the transparency and time requirements. These advantageous factors were empirically established compared to SOTA approaches using the MNIST dataset. Moreover, a list of research questions scrutinizing RL-based solutions, including TRL-HPO, is presented with corresponding recommendations. Future work will target the presented questions to design a general-purpose HPO process.
VIII Acknowledgement
This work was made possible by the facilities of the Shared Hierarchical Academic Research Computing Network (SHARCNET:www.sharcnet.ca) and Compute/Calcul Canada.